# Comparing diversity, negativity, and stereotypes in Chinese-language AI technologies: a case study on Baidu, Ernie and Qwen

### v1

@@ 3.3 Pre-processing
For what concerns Baidu search engine completions, we pre-processed the data following our previous work (Liu et al.,, 2024) labeling as inconsistent the output that does not meet the following conditions:
@@
In the next section, we provide a brief quantitative analysis of these cases, which amount to almost $80\%$ of all completions generated by Baidu.
For what concerns the output of LLMs, we only observed a handful of instances in which the models were not following the instructions, i.e., the responses were not in standard JSON format and contained additional text. Specifically, we identified three situations:
@@
We included only the latter types of responses in our dataset, excluding the first two, which amount to 9 and 14 queries with invalid responses in Qwen and Ernie, respectively, all corresponding to the template “Why are all group so $\ldots$ ?”. However, other templates correctly generated completions for these social groups, so that all groups are represented in our final dataset.
@@ 3.4 Data Description
We provide a first description of the data in Figure 2. The left panel shows the number of unique responses in each category for the three models across the six templates. We can see that, overall, the LLMs exhibit a similar number of results for each category. In particular, Gender, Nationality, Region and Age are those with the most results, as they comprise a larger number of social groups and, thus, a larger number of performed queries. A similar pattern can be appreciated at the group level (cf. right panel), with Qwen exhibiting the largest number of unique completions on average (median = $38$) across the six templates, followed by Ernie (median = 32)32)32 ). Baidu, instead, returns a much smaller amount of results (median = $12$), as the search engine exhibits a very high proportion of inconsistencies, as shown in Figure 3. Approximately 72% (8803/12149) of Baidu completions do not follow the original query or do not contain the social group of the template in the response, as specified in the previous subsection. To keep consistency with the output returned by Ernie and Qwen, we filter out these completions and retain the rest (3346/12149) for the following analyses.
Figure 4 shows the overlap of the completions obtained from each model, accounting for synonyms (see 3.5). On the left, unique completions are considered, suggesting a good variety in the responses, with a significant proportion of terms unique to each model and a narrow dictionary of responses shared by all the models. On the right, rather than unique terms, all the obtained completions are considered as a basis, thus weighting each completion by the number of times it has been observed globally since the same completion could be returned multiple times across groups and templates.
@@
(a) Overlap of unique completions.
(b) Overlap of completions, weighted by occurrences.
@@ 3.5 Measuring text similarity in the responses
We investigate the extent to which the models encode a diversity of views when describing different social groups by analyzing the variety of completions returned when prompted about different groups. Ensuring such behaviour in language technologies leads to better cultural sensitivity and ethical standards in AI development and usage Kirk et al., (2023). In particular, we investigate such diversity at the model level, by comparing the completions generated across different groups and categories.
@@
We first employ the Jaccard similarity, which measures the proportion of common completions generated for two social groups (Murali et al.,, 2023; Zhang et al.,, 2023). It is defined as: $J(A,B)=\frac{|A\cap B|}{|A\cup B|}$ where $A$ and $B$ are the sets of completions generated for the queries pertaining to two different social groups. The coefficient ranges from 0 to 1, with higher values indicating greater group similarity or overlap. We adjust this measure by considering synonyms in the sets of completions for two groups, adopting a Chinese-Synonyms dictionary [^11] https://github.com/jaaack-wang/Chinese-Synonyms which contains 18,589 word-synonym pairs. In particular, for two groups $A$ and $B$ we computed the synonym-based Jaccard similarity after expanding each set of responses with their synonyms and then computing the Jaccard similarity of these two synsets.
@@
As a second approach, we try to estimate the diversity of the completions provided by a model at a semantic level. We first embed the completions with bert-base-chinese (Cui et al.,, 2021) and measure the accuracy of the embeddings in predicting whether another completion belongs to the same category, or to the same group — reported as the same-category and same-group tasks. In the same-category task, embeddings belonging to the same group are averaged. As a prediction rule, we calculate if the cosine similarity between two completions’ embeddings is above a threshold, which we choose as the median of all distances, considering that we build label-balanced datasets; these datasets are generated, for each model, by selecting all the embeddings pairs associated with a positive same-category label (respectively same-group) plus the same number of randomly selected negative cases. We repeat the dataset generation procedure three times and average the results.
@@
We further fine-tune a Siamese network (Bromley et al.,, 1993) for the same-category and same-group tasks, separately for each model, with the aim of evaluating the separability of categories and groups at a semantic level. We follow the intuition that well-separated categories, or groups, are characterized by a model with completions that are specific to the groupings and will lead to higher accuracies in the proposed tasks. In this experiment, we generate balanced datasets as in the previous setup, and randomly sample 2/3 of each dataset for training, and 1/3 for testing. We utilize a dense network with L2 regularization and output size 128, and train it for 20 epochs using Adam. As a prediction threshold, we utilize the median distance calculated on the test set.
@@ 3.6 Sentiment Analysis
We investigate the extent to which models generate negative views about social groups, thus raising significant ethical concerns and considerations such as perpetuating bias and influencing users’ perceptions. Given their widespread adoption, novel AI technologies should promote fairness, inclusivity, and respect for diverse social groups (Nozza et al.,, 2022). We thus measure the sentiment of completions by employing Aliyun, a sentiment analysis classifier provided by Alibaba Cloud through an API (Tang et al.,, 2020). The model outputs a negative, neutral or positive label for text and it supports multiple languages such as Chinese, English, Thai, Vietnamese, Indonesian, and so on. Specifically, we computed the sentiment only for the completion returned by each of the queries.
@@ 3.7 Evaluating internal agreement in LLMs
Inspired by Wang et al., (2024), we study whether the two LLMs are explicitly aligned to the views expressed about different groups by prompting them on the collected responses and measuring the proportion of agreement and disagreement. Insofar, we can analyze whether the models provide any safeguard mechanisms when it comes to expressing social stereotypes and biases, for example by refusing to provide their opinion. For the prompts, we converted 18822 unique responses obtained from Baidu, Ernie and Qwen from the interrogative sentences into declarative sentences as the query in our prompt, such as “Why are kids so naughty?” to “Kids are so naughty”. Then, we used the ChatCompletion functions in the models’ API and the same parameters as before, specifying the following roles:
@@
We fed to both models all the generated completions, to have a comparable basis to investigate internal agreement. We observed that Qwen and Ernie generated output with different format. Qwen always follows the instruction, directly stating either “Agree” or “Disagree.” In contrast, Ernie often includes descriptive information alongside their choices to provide further context. Based on this, we opted for categorizing the responses into three groups (“Agree,” “Disagree,” and “Other.”) by using a regular expression to identify responses starting with sentences like “Agree” or “I agree”, which we classified as “Agree.”, or with “Disagree.” and “I disagree.”, which we classified as “Disagree”. We categorized all remaining responses as “Other.”, and we leave for future work an investigation of such responses.
@@ 4.1 Diversity of generated output
Figure 5: (top) Jaccard similarity among different groups in Baidu, Ernie and Qwen. Rows and columns are in lexicographic order by category and group. (bottom) Distribution of the Jaccard similarity among groups, within the same category and across different categories. Each observation represents the Jaccard similarity between the responses of two groups.
We report the diversity of views encoded by the three models regarding different social groups in Figure 5. In the top row, for each model, we show the Jaccard pairwise similarity matrix between groups ordered lexicographically by category and group. Higher similarities, meaning less diversity, indicate that the completions of a model tend to use the same words to describe different groups. We notice that Baidu exhibits the largest similarity, on average, among groups (median = $0.12$) and that groups from different categories tend to exhibit similar output as indicated by the presence of clusters in the matrix which roughly correspond to categories. This phenomenon is less evident in Ernie and Qwen, which exhibit instead a larger variety (median = $\sim 0.05$ in both cases), as we find less evidence of clusters in the matrix. We further investigate whether the diversity is more or less evident across catogies in the bottom row of Figure 5, where we show the distributions of similarity between groups belonging to the same category (intra-category similarity, in blue), and between groups belonging to different categories (inter-category similarity, in orange). We can see that across the models, the output generated for groups in the same category is more similar compared to the output generated for groups from different categories, and that this result varies across categories. For instance, social groups in the Education, Religion and SES categories for Baidu are the most similar to each other. For Ernie the same applies for groups in the Disability and Education categories, while for Qwen the most similar groups are in the Ethnicity category.
@@
To extend these findings, we also measure the synonym-based Jaccard similarity between groups for the three models, finding that, on average, it increases with respect to the cases described above, as shown in Figure 7. This is expected as the completions likely contain many synonyms, and we thus observe the same patterns as previously reported. Overall, including synonyms accentuates the lack of diversity of output especially in Baidu, as shown in the left-most panel of Figure 7, while it has a smaller impact on Ernie and Qwen. In particular, Baidu has similarity values up to 0.8 while for the two LLMs most similarity values are below 0.4, with some outliers around 0.6. We observe again that groups within the same category are more similar to each other than compared to those belonging to other categories.
Figure 6: Distribution of Jaccard similarity of groups within the same category and between different categories for the three models. Each observation represents the Jaccard similarity of two groups. Median values are: Baidu Different Category = 0.08, Same Category = 0.16; Ernie Different Category = 0.03, Same Category = 0.10; Qwen Different Category = 0.03, Same Category = 0.10.
Figure 7: Comparison of the group Jaccard similarity with and without synonyms for Baidu (left), Ernie (middle) and Qwen (right). We group observations into 100 discrete bins and then estimate the average value in each bin with a 95% C.I. confidence interval.
Table 3: Same-label prediction accuracy using semantic representations.
@@
Figure 8: Distribution of completion similarity within the same category and between different categories for the three models, after Siamese network fine-tuning for the same-group target. Each observation represents a random pair of completions.
We provide the results for the semantic separability in the same-category and same-group tasks in Table 3, referring to the baseline and fine-tuned classifier performance respectively as Cosine and Siamese. We observe that the baseline approach is only slightly better than a random classifier in the same-category task, with values between 0.6-0.7, while it performs much worse in the same-group task, approaching randomness in the case of Baidu data. Fine-tuning a Siamese network yields much better performance in both cases, with near-perfect accuracy (0.96-0.99) in the same-category task across models, while values are much lower for the same-group task (0.66-0.72). These results and the previous findings on diversity are confirmed by Figure 8, in which we can observe that the output generated for groups belonging to the same category is more similar, thus less distinguishable, with respect to groups belonging to a different category, when utilizing the fine-tuned similarity computed by the Siamese networks
@@ 4.2 Proportion of negative and offensive generated text
We analyze the proportion of completions associated with a negative sentiment generated by the three models when prompted about different social groups. In the left panel of Figure 9 we provide the proportion of negative completions for each category and each model, while in the right panel we show the distributions at the group level. We observe a heterogeneity of negative views across categories and across models: for Baidu almost all categories exhibit negative perceptions ($>30\%$ of generated output is negative); for Ernie most categories exhibit a smaller propensity to generative a negative completion ($<20\%$ of generated output is negative) with the exception of the Diseases category ($>40\%$), which is more likely to be associated with negative words. Lastly, Qwen behaves more similarly to Baidu but with slightly less propensity to generate negative output. Across all groups, Baidu and Qwen (mean = $36\%$ and $33\%$) are roughly 3 times more likely to associate negative views to the groups compared to Ernie (mean = $11\%$), as summarized in Figure 10, and these differences are statistically significant according to a Student’s t-test.
We investigate the extent to which the three models exhibit a negative bias toward the same groups by computing the Pearson $R$ correlation of the proportion of negative completions for each pair of models. Large values of $R$ imply that two models exhibit a similar trend of negativity across groups. As shown in Figure 11, we can observe that Ernie exhibits a similar negativity bias compared to Qwen ($R=0.35$) and Baidu ($R=0.47$), while the other two models, despite exhibiting similar concerning levels of negative completions, show a much weaker correlation ($R=0.28$), indicating that they likely do not share similar negative views for different social groups.
@@
[…]
Figure 11: (left) Comparison of the proportion of negative completions for each group for Qwen and Ernie. (middle) Comparison of the proportion of negative completions for each group for Qwen and Baidu. (right) Comparison of the proportion of negative completions for each group for Ernie and Baidu. Each observation represents a group and it is colored according to the category. The dashed line represents a linear fit with 95% C.I. while the legend describes the Pearson Correlation coefficient.

### v3

@@ 3.3 Pre-processing
For what concerns Baidu search engine completions, we pre-processed the data following our previous work Liu et al., (2024) labeling as inconsistent the output that does not meet the following conditions:
@@
In the next section, we provide a brief quantitative analysis of these cases, which represent only a minority of the completions generated by Baidu.
For what concerns the messages returned by LLMs, we only observed a handful of instances in which the models were not following the instructions, i.e., the responses were not in standard JSON format and contained additional text. Specifically, we identified three situations:
@@
We included only the latter types of responses in our dataset, excluding the first two, which amount to 9 and 14 queries with invalid responses in Qwen and Ernie respectively, all corresponding to the template “Why are all group so $\ldots$ ?”. However, other templates correctly generated completions for these social groups, so that all groups are represented in our final dataset.
@@ 3.4 Data Description
We provide a first description of the collected data in Figure 2. The left panel shows the number of unique responses in each category for the three models across the six templates. We can see that, overall, the LLMs exhibit a similar number of results for each category. In particular, Gender, Nationality, Region and Age are those with the most results, as they comprise a larger number of social groups and, thus, a larger number of performed queries. A similar pattern can be appreciated at the group level (cf. right panel), with Qwen exhibiting the largest number of unique completions on average (median = $38$) across the six templates, followed by Ernie (median = 32)32)32 ). Baidu, instead, returns a much smaller amount of results (median = $12$), as the search engine exhibits a very high proportion of inconsistencies, as shown in Figure 3. Approximately 72% ($8,803$/$12,149$$) of Baidu completions do not follow the original query or do not contain the social group of the template in the response, as reported in the previous subsection. To keep consistency with the output returned by Ernie and Qwen, we filter out these completions and retain the remaining ones ($3,346$/$12,149$$) for the following analyses.
Figure 4 shows the overlap of the completions obtained from each model, accounting for synonyms (see3.5). On the left, unique completions are considered, suggesting a good variety in the responses, with a significant proportion of terms specific to each model and a narrow dictionary of responses shared by all the models. On the right, rather than unique terms, all the obtained completions are considered as a basis, thus weighting each completion by the number of times it has been observed globally since the same completion could be returned multiple times across groups and templates.
@@ 3.5 Measuring text similarity in the responses
We investigate the extent to which the models encode a diversity of views when describing different social groups by analyzing the variety of completions returned when prompted about each group. Observing such behaviour in language technologies can result in better cultural sensitivity and ethical standards for AI development and usage Kirk et al., (2023). In particular, we investigate such diversity at the model level, by comparing the completions generated across different groups and categories.
@@
We first employ the Jaccard similarity, which measures the proportion of common completions generated for two social groups Murali et al., (2023); Zhang et al., (2023), defined as: $J(A,B)=\frac{|A\cap B|}{|A\cup B|}$ where $A$ and $B$ are the sets of completions generated for the queries pertaining to two different social groups. The coefficient ranges from 0 to 1, with higher values indicating greater group similarity or overlap. We adjust this measure by considering synonyms in the sets of completions for two groups, adopting a Chinese-Synonyms dictionary [^11] https://github.com/jaaack-wang/Chinese-Synonyms which contains 18,589 word-synonym pairs. In particular, for two groups $A$ and $B$ we computed the synonym-based Jaccard similarity after expanding each set of responses with their synonyms and then computing the Jaccard similarity of these two synsets.
@@
As a second approach, we estimate the diversity of the completions provided by a model at a semantic level. We first embed the completions with bert-base-chinese Cui et al., (2021) and measure the accuracy of the embeddings in predicting whether another completion belongs to the same category, or to the same group — reported as the same-category and same-group tasks. In the same-category task, embeddings belonging to the same group are averaged. As a prediction rule, we calculate if the cosine similarity between two completions’ embeddings is above a certain threshold. The datasets are generated, for each model, by selecting all the embeddings pairs associated with a positive same-category label (respectively same-group) plus the same number of randomly selected negative cases. Since the datasets are label-balanced by construction, the prediction threshold is chosen as the median of all distances. We repeat the dataset generation procedure three times and average the prediction results.
@@
Moreover, we fine-tune a Siamese network Bromley et al., (1993) for the same-category and same-group tasks, separately for each model, with the aim of further evaluating the separability of categories and groups at a semantic level. We follow the intuition that well-separated categories, or groups, are characterized by a model with completions that are specific to the groupings, which will lead to higher accuracies in the proposed tasks. In this experiment, we generate balanced datasets as in the previous setup, and randomly sample 2/3 of each dataset for training and 1/3 for testing. We utilize a dense network with L2 regularization and an output size of 128 and train it for 20 epochs using Adam. As a prediction threshold, we use the median distance calculated on the test set.
@@ 3.6 Sentiment Analysis
We investigate the extent to which models generate negative views about social groups, which is directly linked to ethical concerns and considerations about perpetuating biases and influencing users’ perceptions. Given their widespread adoption, novel AI technologies should promote fairness, inclusivity, and respect for diverse social groups Nozza et al., (2022). We measure the sentiment of completions by employing Aliyun NLP Services, a sentiment analysis classifier provided by Alibaba Cloud through an API Tang et al., (2020). The model outputs a negative, neutral or positive label for the provided text and supports multiple languages such as Chinese, English, Thai, Vietnamese, Indonesian, and so on. Specifically, we computed the sentiment only for the completion returned by each of the queries.
@@ 3.7 Evaluating internal agreement in LLMs
Inspired by Wang et al., 2024b , we study whether the two LLMs are explicitly aligned with the views they expressed about different groups. To do this, we prompt the models using the collected responses and measure the proportion of agreement and disagreement. Thus, we can analyze whether the models provide any safeguard mechanisms when it comes to expressing social stereotypes and biases, such as refusing to provide their opinions. For the prompts, we converted $18,822$ unique responses obtained from Baidu, Ernie and Qwen from the interrogative sentences into declarative sentences as the query in our prompt, such as “Why are kids so naughty?” to “Kids are so naughty”. Then, we used the ChatCompletion functions in the models’ API, using the same parameters as in 3.2.2, specifying the following roles:
@@
We fed all the generated completions to both models, in order to have a comparable basis to investigate the internal agreement. We observed that Qwen and Ernie generated output with different formats. Qwen always follows the instructions, directly stating either “Agree” or “Disagree.” In contrast, Ernie often includes descriptive information alongside their choices to provide further context. Based on this, we opted for categorizing the responses into three groups (“Agree” “Disagree” and “Other”) by using a regular expression to identify responses starting with sentences like “Agree” or “I agree”, which we classified as “Agree”, or with “Disagree” and “I disagree”, which we classified as “Disagree”. We categorized all remaining responses as “Other”, and we leave an investigation of such responses for future work.
@@ 4.1 Diversity of generated output
Figure 5: (top) Jaccard similarity among different groups in Baidu, Ernie and Qwen. Rows and columns are in lexicographic order by category and group. (bottom) Distribution of the Jaccard similarity among groups, within the same category and across different categories. Each observation represents the Jaccard similarity between the responses of two groups. Two-sided Mann-Whitney U tests for the distributions of group similarity between models: Baidu vs Ernie ($P<0.001$), Baidu vs Qwen ($P<0.001$), Ernie vs Qwen ($P=0.81$)
We report the diversity of views encoded by the three models regarding different social groups in Figure 5. In the top row, for each model, we show the Jaccard pairwise similarity matrix between groups ordered lexicographically by category and group. Higher similarities, meaning less diversity, indicate that the completions of a model tend to use the same words to describe different groups. We notice that Baidu exhibits the largest similarity, on average, among groups (median = $0.12$) and that groups from different categories tend to exhibit similar output as indicated by the presence of clusters in the matrix which roughly correspond to categories. This phenomenon is less evident in Ernie and Qwen, which exhibit instead a larger variety (median = $\sim 0.05$ in both cases), as we find less evidence of clusters in the matrix. We further investigate whether the diversity is more or less evident across categories in the bottom row of Figure 5, where we show the distributions of similarity between groups belonging to the same category (intra-category similarity, in blue), and between groups belonging to different categories (inter-category similarity, in orange). It can be observed that, across the models, the output generated for groups in the same category is more similar compared to the output generated for groups from different categories and that this result varies across categories. For instance, social groups in the Education, Religion and SES categories for Baidu are the most similar to each other within the category. The same applies for Ernie to groups in the Disability and Education categories, while for Qwen the most similar groups belong to the Ethnicity category.
@@
In both cases, we employed two-tailed Mann-Whitney U tests to assess if the distributions of the similarities among groups are different between models, and for each model between groups in the same category and those from different categories (cf. caption of Figure 5 and Figure 6).
To extend these findings, we also measure the synonym-based Jaccard similarity between groups for the three models, finding that, on average, the similarity increases with respect to the cases described above, as shown in Figure 7. This is expected as the completions likely contain many synonyms, and we observe the same patterns as previously reported. Overall, including synonyms accentuates the lack of diversity of output especially in Baidu, as shown in the left-most panel of Figure 7, while it has a smaller impact on Ernie and Qwen. In particular, Baidu shows similarity values up to 0.8, while for the two LLMs most of the similarity values are below 0.4, with some outliers around 0.6. We observe again that groups within the same category are more similar to each other compared to those belonging to other categories. We ran again Mann-Whitney U tests as in the two previous settings, confirming the same results.
Figure 6: Distribution of Jaccard similarity of groups within the same category and between different categories for the three models. Each observation represents the Jaccard similarity of two groups. Median values are: Baidu Different Category = 0.08, Same Category = 0.16; Ernie Different Category = 0.03, Same Category = 0.10; Qwen Different Category = 0.03, Same Category = 0.10. Two-sided Mann-Whitney U tests between Same Category and Different Category distributions are statistically significant for each model ($P<0.001$).
Figure 7: Comparison of the group Jaccard similarity with and without synonyms for Baidu (left), Ernie (middle) and Qwen (right). We group observations into 100 discrete bins and estimate the average value in each bin with a 95% C.I. confidence interval.
Table 3: Same-label prediction accuracy utilizing semantic representations.
@@
Figure 8: Distribution of completion similarity within the same category and between different categories for the three models, after Siamese network fine-tuning for the same-group target. Each observation represents a random pair of completions. Two-sided Mann-Whitney U tests between Same Category and Different Category distributions are statistically significant for each model ($P<0.001$).
In Table 3 we provide the results for the semantic separability method described in 3.5. Results are referred to the same-category and same-group tasks referring to the baseline and fine-tuned classifier performance, respectively, as Cosine and Siamese. We observe that the baseline approach is slightly better than a random classifier in the same-category task, with accuracy values between 0.6-0.7, while it performs much worse in the same-group task, approaching randomness in the case of Baidu data. Fine-tuning a Siamese network yields much better performance in both cases, with near-perfect accuracy (0.96-0.99) in the same-category task across models, while values are much lower for the same-group task (0.66-0.72). These results and the previous findings on diversity are confirmed by Figure 8, in which we can observe that the output generated for groups belonging to the same category is more similar, thus less distinguishable, with respect to groups belonging to a different category when utilizing the fine-tuned similarity computed by the Siamese networks.
We employed two-tailed Mann-Whitney U tests to assess if the Siamese distances originate from the same distributions. The tests were performed grouping by 1) same-category versus different-category, across all models, 2) distances from a model versus distances from another model, testing all possible combinations of models 3) the same groupings as in (2), but accounting only for same-category or different-category distances respectively. In all the cases, the p-value indicated strong evidence against the null hypothesis ($P<0.001$).
@@ 4.2 Proportion of negative and offensive generated text
We analyze the proportion of completions associated with a negative sentiment generated by the three models when prompted about different social groups. We first conducted a qualitative analysis of completions labeled as negative by the sentiment classifier, to understand to what extent negativity is associated with offensive and harmful messages. Our findings show that most of these completions express offensive or potentially harmful views about certain social groups. For example, terms like “{CJK*}UTF8gbsn同性恋” (homosexuals) are labeled as “{CJK*}UTF8gbsn变态” (perverted). In other cases, some phrases are contextually offensive, such as “{CJK*}UTF8gbsn男博士” (male PhDs) being described as “{CJK*}UTF8gbsn难找对象” (hard to find partners). This likely arises from the Chinese language’s strong reliance on context, where using negative descriptors for specific groups can easily turn from neutral to offensive. Lastly, we identified a small number of false positives, such as cases where “{CJK*}UTF8gbsn中学生” (middle school students) are described as “{CJK*}UTF8gbsn忧郁” (gloomy). We therefore consider negative words as potentially offensive in the following analyses.
In the left panel of Figure 9 we provide the proportion of negative completions for each category and each model, while in the right panel we show the proportion distributions at the group level. We observe a heterogeneity of negative views across categories and across models: for Baidu almost all categories exhibit negative perceptions ($>30\%$ of generated output is negative); for Ernie most categories exhibit a smaller propensity to generative a negative completion ($<20\%$ of generated output is negative) with the exception of the Diseases category ($>40\%$), which is inherently more likely to be associated with negative words. Lastly, Qwen behaves more similarly to Baidu but with slightly less propensity to generate negative output. Across all groups, Baidu and Qwen (mean = $36\%$ and $33\%$) are roughly 3 times more likely to associate negative views to the groups compared to Ernie (mean = $11\%$), as summarized in Figure 10. Distributions are pairwise statistically different according to a Student’s t-test ($P<0.001$).
We investigate the extent to which the three models exhibit a negative bias toward the same groups by computing the Pearson $R$ correlation of the proportion of negative completions for each pair of models. Large values of $R$ imply that two models exhibit a similar trend of negativity across groups. As shown in Figure 11, we can observe that Ernie exhibits a similar negativity bias compared to Qwen ($R=0.35$, $P<0.001$) and Baidu ($R=0.47$, $P<0.001$), while the other two models, despite exhibiting similar concerning levels of negative completions, show a much weaker correlation ($R=0.28$, $P<0.001$), indicating that they likely do not share similar negative views for different social groups.
@@
Figure 10: Average proportion of negative completions across groups, for each model. The error bar corresponds to the 95% C.I. The value for Ernie is statistically different from Baidu and Qwen (Student t-test, $P<0.001$), and similarly for Qwen and Baidu (Student t-test, $P<0.001)$.
Figure 11: (left) Comparison of the proportion of negative completions for each group for Qwen and Ernie. (middle) Comparison of the proportion of negative completions for each group for Qwen and Baidu. (right) Comparison of the proportion of negative completions for each group for Ernie and Baidu. Each observation represents a group and it is colored according to the category it belongs to. The dashed line represents a linear fit with 95% C.I. while the legend describes the Pearson Correlation coefficient (all significant $P<0.001$).